Displacement of a particle is given by X equal to 20 + 10t² square calculate the instantaneous velocity t=2s

Grade:12

1 Answers

Pari
3 years ago
we are given that
X = 20 + 10t2 , x is the position of the object and t is the time.
by Diffrenetiation , we will get the velocity i.e. dx/dt = V
dx/dt of 20 + 10t2 = 20t
now T= 4seconds (given in the qustion)
20 x 2 += 40m/s
answer= 40m/s at t= 2 seconds